OpenStudy (anonymous):

At noon on a clear day, sunlight reaches the earth\'s surface at Madison, Wisconsin, with an average power of approximately 3.00 kJ·s–1·m–2. If the sunlight consists of photons with an average wavelength of 510.0 nm, how many photons strike a 2.70 cm2 area per second?

OpenStudy (anonymous):

An intensity of 3.00 kJ m^-2 gives:\[3000/10000*2.7=0.81\]0.81 J per 2.70 cm^2 per second. Each photon has an energy of:\[hc/\lambda=3.89499154*10^{-19} joules\] So the total number of photons:\[0.81/(3.895*10^{-19})=2.08 * 10^{18}\]